if y=(sin ?¹x)+(sin ?¹(√1−x²))then find dy÷dx.

Grade:12

1 Answers

Parvez ali
10 years ago
y=sin-1 x +sin -1root(1-x^2)
=sin-1x +cos-1x=pie/2
dy/dx=0
Parvez Ali
Askiitian Faculty
B-tech
ISM Dhanbad